what are five points that i can graph for the equation y=x^2+4 (vertex form)
and for the equation y=2x^2+6x=12 (standard form)

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

what are five points that i can graph for the equation y=x^2+4 (vertex form)
------------
Pick 5 values for x and find y.
---------------------------------
and for the equation y=2x^2+6x=12 (standard form)
Too many equal signs in that one.

Answer by Edwin McCravy(20060)

what are five points that i can graph for the equation y=x^2+4 (vertex form)
The vertex form is

y = 1(x-0)²+4
y = a(x-h)²+k, so the vertex is (h,k) = (0.4)

(-3,13), (-2,8), (-1,5), (0,4), (1,5), (2,8), (3,13)

----------------------

and for the equation

y=2x^2+6x=12
You mistyped an = sign for either a + or a - sign

If you meant a + sign, y=2x²+6x+12

(-4,20), (-3,12), (-2,8), (-1.5,7.5), (-1,-8), (0,12), (1,20)

If you meant a - sign, y=2x²+6x-12 

(-5,8), (-4,-4), (-3,-12), (-2,-16), (-1.5,-16.5), (-1,-16), 
(0,-12), (1,-4), (2,8)

Edwin